chapter 5 probability

0.0(0)
Studied by 0 people
call kaiCall Kai
learnLearn
examPractice Test
spaced repetitionSpaced Repetition
heart puzzleMatch
flashcardsFlashcards
GameKnowt Play
Card Sorting

1/16

encourage image

There's no tags or description

Looks like no tags are added yet.

17 Terms

1
New cards

random process

generates outcomes determined purely by chance

2
New cards

trial

one repetition of a random process

3
New cards

probability

chance process is a number between 0-1 that describes the proportion of times the outcome would occur in series of repetitions

4
New cards

law of large numbers

observing more & more repetitions of any chance process the proportion of times that a specific outcome occurs approaches its probability

5
New cards

probability model

description of some chance process that consists of two parts

  • all possible outcomes

  • probability for each outcome


6
New cards

sample space

list of all possible outcomes

7
New cards

mutually exclusive events

events a & b are mutually exclusive if it is impossible for the events to occur at the exact same time from single trial — P (A and B) = 0

8
New cards

addition rule

P (A or B) = P(A) + P(B) - P(A and B)

9
New cards

complement

A^c — contains exactly the outcomes not in A

10
New cards

intersection

(A & B) — the set of all outcomes in both events

11
New cards

union

(A or B) — set of outcomes in either event

12
New cards

conditional probability

probability that one event happens given another event is known to have happened

  • not possible for events to be independent & mutually exclusive


13
New cards

independent events

events A & B are independent if one of them occurring does not change the probability of the other event occurring

14
New cards

tree diagram

shows the sample space of a chance process involving multiple stages

  • probability of each outcome is show on corresponding branch

  • after first stage the probabilities are conditional


15
New cards

probability of at least one

P (at least one) = 1- P(event never occurs) ^n)

  • n = how many don’t


16
New cards

multiplication rule (dependent events)

P (A and B) = P(A) x P (B given A)

17
New cards

multiplication rule (independent events)

P (A & B) = P(A) x P(B)